OpenStudy (anonymous):

Can some one please explain to me how to do trigonometric integrations like the integral of (4 cos(x)/(1+sin(x))?

OpenStudy (kirbykirby):

Let \(u = 1+\sin x\) \(du = \cos x \, dx \) :)

OpenStudy (kirbykirby):

\[\large \int\frac{4\cos x}{\sin x}\, dx \] If \(u = 1 + sin x\) \(du = \cos x \, dx\) \[ \large \int\frac{4}{u}\, du\] and recall that: \[\int \frac{1}{x}\, dx=\ln|x|+C \]

OpenStudy (anonymous):

thank you i think i understand it now (:

OpenStudy (kirbykirby):

:)